Zero Turbulence Airline provides air transportation services between Los Angeles, California, and Kona, Hawaii. A single Los Angeles to Kona round-trip flight has the following operating statistics:

Fuel $7,000

Flight crew salaries 3,200

Airplane depreciation 3,480

Variable cost per passenger—business class 140

Variable cost per passenger—economy class 120

Round-trip ticket price—business class 800

Round-trip ticket price—economy class 300

It is assumed that the fuel, crew salaries, and airplane depreciation are fixed, regardless of the number of seats sold for the round-trip flight.

a. Compute the break-even number of seats sold on a single round-trip flight for the overall product, E. Assume that the overall product mix is 10% business class and 90% economy class tickets.

Total number of seats at break-even seats ____???

b. How many business class and economy class seats would be sold at the break-even point?

Business class seats at break-even seats ____???

Economy class seats at break-even seats ____???
Answer #1

FIXED COST = Fuel + Flight crew salaries +Airplane depreciation

= 7000 + 3200 + 3480

= 13680

Contribution = round trip ticket price – variable cost per passenger

Business class= 800 – 140 = 660

Economy class = 300 – 120 = 180

Product mix = 10% business class and 90% economy class tickets

Overall contribution= 660 *0.10 + 180 * 0.90 = 228

Contribution margin = 228/350*100 = 65

Total number of seats at break-even seats = 13680/228 = 60

b. Business class and economy class seats would be sold at the break-even point

Business class seats at break-even seats = 60* 0.10 = 6

Economy class seats at break-even seats = 60* 0.9 = 54